What should be the energy of M orbit of hydrogen atom if the energy of its first orbit is -864ev?

Grade:11

1 Answers

Vikas TU
7 years ago
Energy at first orbit = k*z^2/n^2 (k = constant)
                            = k* 1/1 = -864 ev (given)..........................(A)
 
Energy at Mth orbit = k* 1/M^2 = E (let)..........................(B)

Divide (A) and (B) eqns.........
-864/E = M^2
E = -864/M^2 ev.